1. During a downhill ski race, the distance (in yds) a skier has travelled since leaving the starting gate follows this quadratic equation: d=5t^2+32t+1000 where t is the time in seconds measuring her run. After how many seconds has she skied 2,000yds?
I have tried this question and keep getting stuck. please help

Answer by rothauserc(4718)

we are given d=5t^2+32t+1000 and the skier has skied 2000 yards
2000 = 5t^2+32t+1000 and solve for t
5t^2+32t-1000 = 0
use quadratic formula
t = (-b + square root(b^2 - 4ac)) / 2a = (-32 + square root(32^2 + 20*1000) / 10 = 11.30 seconds
t = (-b - square root(b^2 - 4ac)) / 2a = (-32 - square root(32^2 + 20*1000) / 10 = -17.70 seconds
our answer is t = 11.30 seconds